OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, 3) and perpendicular to y = -1/3 x + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@TrapBoiTrynaEducaate

OpenStudy (anonymous):

@princeharryyy

OpenStudy (amistre64):

well, we know that 3x is going to have to be a part of it right?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (amistre64):

choose the option that when x=-1, y=3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

D ?

OpenStudy (amistre64):

lets test that option ... 3 = 3(-1) is this a true statement?

OpenStudy (amistre64):

one idea, if trial and error is not your strong point is to solve the setup 3 = 3(-1) + b to find the correct ending term

OpenStudy (princeharryyy):

y = 3x + 6

OpenStudy (anonymous):

but its not a true statement

OpenStudy (princeharryyy):

b is correct @IndianBoyNiko

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

The equation of a line is shown below: y = - x - 7 What is the equation of a line which is parallel to this line and passes through (1, 79)? y = -x + 80 y = x - 80 y = 7x - 80 y = -7x + 80

OpenStudy (anonymous):

@princeharryyy

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

It must have the same slope so you can already take out 2 options.

OpenStudy (anonymous):

which ones ?

OpenStudy (sleepyjess):

Actually you can take 3 out

OpenStudy (anonymous):

which ones ?

OpenStudy (sleepyjess):

The first term MUST be -x

OpenStudy (anonymous):

so A ?

OpenStudy (princeharryyy):

y = -x + 80

OpenStudy (sleepyjess):

Yes and to confirm that substitute 1 for x

OpenStudy (princeharryyy):

@IndianBoyNiko

OpenStudy (anonymous):

Thank you

OpenStudy (sleepyjess):

@princeharryyy Do not give out direct answers please.

OpenStudy (anonymous):

The first five terms of a sequence are shown below. 5, 9, 13, 17, 21 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 5n - 1; n ≥ 1 f(n) = 5n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1

OpenStudy (anonymous):

This one was the hardest but I got C

OpenStudy (anonymous):

but idk if it's right

OpenStudy (sleepyjess):

One question per post please

OpenStudy (princeharryyy):

f(n) = 4n + 1; n ≥ 1 simply plugin and check verifies all

OpenStudy (sleepyjess):

Substitute in a number. f(1)= 5(1) - 1; n ≥ 1 f(n) = 5(1) + 1; n ≥ 1 f(n) = 4(1) - 1; n ≥ 1 f(n) = 4(1) + 1; n ≥ 1